OpenStudy (anonymous):

x^2-6x+y^2+14y+33=0 can someone please explain to me how to get the radius and center of this. i am so confused

OpenStudy (anonymous):

you have to complete the square

OpenStudy (anonymous):

Complete the square on : x^2-6x Complete the square on y^2+14y

OpenStudy (anonymous):

x^2-6x+____+y^2+14y+____=-33 x^2-6x+(1/2(6))^2+y^2+14y+(1/2(14))^2=-33+9+49 (x-3)^2+(y+7)^2=r^2 which is -33+58 which on a good day to me is 25 so the equation looks like this \[(x-3)^2+(y+7)^2\]

OpenStudy (anonymous):

\[(x-3)^2+(y+7)^2=25\]

OpenStudy (anonymous):

the reason i added 9 and 49 to -33 is because thats the "completing the square" that you see on the left (1/2(6))^2=9 and (1/2(14))^2=49 and you have to add it to the other side because its as if your adding "0" to the problem, so your allowed to do so.

OpenStudy (anonymous):

then the center is at (k,h)= (x-k)^2+(y-h)^2=r^2 center of your problem is (3,-7) with a radius at 5 is this making sense?